# Languages of Boundedly-Ambiguous Vector Addition Systems with States

### v1

@@ Abstract
TODO
@@
TODO
@@ 1 Introduction
Determinism is a central notion in computer science. Deterministic systems often allow for more efficient algorithms. On the other hand, usually deterministic systems are less expressive, so in many cases there is no equivalent deterministic system and cannot use the more efficient techniques. For those reasons, there is recently a lot of research devoted to various notions restricting nondeterminism in a milder way than determinism. The hope is that systems having the considered properties are more expressive than the deterministic ones, but still allow for robust algorithms design. One prominent example of such a notion is unambiguity; a system is unambiguous if for every word there is at most one accepting run over this word. In the last decade unambiguous systems were intensely studied and for various classes of infinite-state systems the unambiguous case turns out to be much more tractable [DBLP:conf/dcfs/Colcombet15, DBLP:conf/icalp/Raskin18, DBLP:conf/stacs/MottetQ19, DBLP:conf/concur/CzerwinskiFH20, DBLP:conf/icalp/GoosK022]. Similar notions were also investigated recently, like $k$-ambiguity (each word is accepted by at most $k$ runs) and history-determinism (a weakened version of determinism), in both cases one can design more efficient algorithms in some cases [DBLP:conf/concur/CzerwinskiH22, history-deterministic].
@@
In this paper we focus on studying milder version of determinism for Vector Addition Systems with States (VASS). VASSs and related Petri nets are popular and fundamental models of concurrency with many applications both in theory and in practical modelling [DBLP:journals/siglog/Schmitz16, Section 5]. Languages of VASSs with restricted nondeterminism were already studied for several years, mostly with the acceptance condition being the set of accepting states. In [DBLP:conf/concur/CzerwinskiFH20] it was shown that the universality problem for unambiguous VASSs is decidable in ExpSpace, in contrast to Ackermann-completeness of the problem for VASSs without that restriction. In [DBLP:conf/concur/CzerwinskiH22, DBLP:journals/lmcs/CzerwinskiH25] the language equivalence problem was considered for unambiguous VASS and more generally for $k$-ambiguous VASSs and it was shown to be decidable and Ackermann-complete, in contrast to undecidability in general [DBLP:journals/tcs/ArakiK76], even in dimension one [DBLP:conf/lics/HofmanMT13, Thm. 20]. The choice of universality and equivalence problems is deliberate here. The complexity of the seemingly more natural language emptiness problem (or equivalently of the reachability problem) does not depend on the unambiguity assumption. Indeed, one can always transform a system to a deterministic one by assigning all the transitions unique labels without affecting nonemptiness. Therefore, in order to observe a difference in complexity or decidability after restricting nondeterminism, one should consider problems, which do not simply ask about existence of some object. Some natural problems concerning languages of that type are the universality problem and the language equivalence problem and these two problems were extensively studied for many models with the unambiguity assumption or some modification of it. First of all, the universality and equivalence problems are solvable in PTime [DBLP:journals/siamcomp/StearnsH85] (and even in NC2 [DBLP:journals/ipl/Tzeng96]) for unambiguous finite automata (UFA), in contrast to PSpace-hardness of both problems for NFA. The universality problem was also shown decidable for unambiguous register automata in a sequence of papers [DBLP:conf/stacs/MottetQ19, DBLP:conf/stacs/BarloyC21, DBLP:conf/icalp/CzerwinskiMQ21] with improving complexity, which contrasts undecidability in the case without restricted nondeterminism [DBLP:journals/tocl/NevenSV04, Thm 5.1]. This line of research culminated in a very elegant contribution [DBLP:conf/lics/BojanczykKM21, DBLP:journals/theoretics/BojanczykFKM24], which showed in particular that not only the universality problem, but also the equivalence problem is solvable in ExpTime for unambiguous register automata, and in PTime in the case of fixed number of registers. Another popular restriction of nondeterminism was also studied recently: history-determinism. A system is history-deterministic if its nondeterminism can be resolved on the fly, based on the history of a particular run. The equivalence and inclusion problems were shown to be decidable for one-dimensional history-deterministic VASSs [DBLP:conf/fossacs/PrakashT23], but undecidable for two-dimensional history-deterministic VASSs [history-deterministic].
@@
Despite a lot of research on algorithms for unambiguous and boundedly-ambiguous ($k$-ambiguous for some $k\in\mathbb{N}$) VASS not much is known about the class of languages recognisable by such models. In particular, to our best knowledge, till now it was even not known whether there exist any VASS language (we consider acceptance by states), which is not a language of an unambiguous VASS. The reason behind this lack of knowledge was absence of any technique, which can show that a given language of a VASS cannot be recognised by an unambiguous VASS. The quest for such a technique is natural and the question deserves investigation. Analogous problems were considered for other models of computation. For finite automata the question trivialises, as deterministic automata recognise all the regular languages. However, already for weighted automata over a field the problems are highly nontrivial and were recently studied in-depth [DBLP:conf/lics/BellS23, DBLP:journals/corr/abs-2410-03444], in particular it is decidable whether given weighted automaton is unambiguisable, so equivalent to some unambiguous weighted automaton [DBLP:conf/lics/BellS23]. The problem whether a given context-free language is unambiguisable is known to be undecidable since the 60-ties [undecidable_cfl, DBLP:journals/mst/Greibach68]. The aim of this paper is deepening the understanding of the class of languages recognisable by boundedly-ambiguous VASS and its subclasses.
@@
For any class $\mathcal{C}$ of languages containing all the regular languages and contained in the class of all boundedly-ambiguous VASS languages it is undecidable to check whether language of a given $1$-VASS accepting by states belongs to $\mathcal{C}$.
@@
It is undecidable whether language of a given $1$-VASS accepting by states is recognisable by some
@@
  • $k$-ambiguous $1$-VASS for given $k\in\mathbb{N}$.
@@
Organisation of the paper. In Section 2 we introduce preliminary notions and recall useful lemmas. Section 3 is devoted to showing the two main technical tools, namely Lemmas 3.2 and 3.23. In Section 4 we present closure properties and expressivity results for the classes of $k$-ambiguous VASSs. Next, in Section 5 we show our main result, namely Theorem 1.1. Theorem 6.1 is proved in Section 6. Finally, in Section 7 we discuss interesting future research directions.
@@ 2 Preliminaries
Basic notions. For $a,b\in\mathbb{N}$ such that $a\leq b$ we write $[a,b]$ for the set of integers $\{a,a+1,\ldots,b\}$. For $a\in\mathbb{N}$ we write $[a]$ for the set $[1,a]$. For a vector $v\in\mathbb{Z}^{d}$ we write $v_{i}$ for the $i$-th entry of $v$. For a vector $v\in\mathbb{Z}^{d}$ we write $\textup{support}(v)$ for the set $\{i\mid v_{i}>0\}$. For two vectors $v,u\in\mathbb{N}^{d}$ we write $v\geq u$ if for all $i\in[d]$ we have $v_{i}\geq u_{i}$. We also extend this order to ($(\mathbb{N}\cup\{\omega\})^{d}$ where $\omega$ is bigger than any natural number. We write $\mathbb{N}_{\omega}$ for the set $\mathbb{N}\cup\{\omega\}$. Whenever we speak about the norm of vector $v\in\mathbb{Z}^{d}$ we mean $||v||=\max_{1\leq i\leq d}|v_{i}|$. For a word $w$ we denote by #$\#_{a}(w)$ the number of letters $a$ in the word $w$. Downward-closed sets. For two vectors $u,v\in\mathbb{N}^{d}$ we say that $u\preceq v$ if for all $i\in[1,d]$ we have $u_{i}\leq v_{i}$. A set $S\subseteq N^{d}$ is downward-closed if for each $u,v\in\mathbb{N}^{d}$ it holds that $u\in S$ and $v\preceq u$ implies $v\in S$. For $u\in\mathbb{N}^{d}$ we write $u\downarrow$ for the set $\{v\mid v\preceq u\}$. If a downward-closed set is of the form $u\downarrow$ we call it a down-atom. Observe, that a one-dimensional set $S\subseteq\mathbb{N}$ is downward-closed if either $S=\mathbb{N}$ or $S=[0,n]$ for some $n\in\mathbb{N}$. Thus we have either $S=\omega\downarrow$ in the first case or $S=n\downarrow$ in the second case. We call a downward-closed set $D\subseteq\mathbb{N}^{d}$ a down-atom if it is of a form $D=D_{1}\times\ldots\times D_{d}$, where for all $i\in[1,d]$ we have that $D_{i}$ is a downward-closed one dimensional set. For simplicity we write $D=(n_{1},n_{2},\ldots,n_{d})\downarrow$ if $D=n_{1}\downarrow\times n_{2}\downarrow\ldots n_{d}\downarrow$. Therefore each down-atom is of the form $u\downarrow$ where $u\in(\mathbb{N}\cup\omega)^{d}$. The following proposition will be helpful in our considerations.
Proposition 2.1 ([CzerwinskiLMMKS18] Lemma 17, [Dickson]).
@@
Vector Addition Systems with States. A $k$-dimensional Vector Addition System with States ($k$-VASS) is a nondeterministic finite automaton with $k$ non-negative integer counters. Transitions of the VASS manipulate these counters. Formally, we can define VASS $V$ as $V=(\Sigma,Q,\delta,I,F)$ where $\Sigma$ is a finite alphabet, $Q$ is a finite set of automaton states, $\delta$ is a transition relation $\delta\subseteq Q\times(\Sigma\cup\varepsilon)\times\mathbb{Z}^{k}\times Q$, $c_{0}\in Q\times\mathbb{N}^{k}$ is an initial configuration, $I$ is a finite set of initial configurations and $F\subseteq Q\times\mathbb{N}^{k}$ is a finite set of final configurations. For a transition $t=(s,a,v,s^{\prime})\in\delta$ we say, that transition is over $a$ or the transition reads letter $a$. We also write $\textsc{eff}(t)$ for the effect of transition, which is $v$. We define the norm of transition $t$ as the norm of $v$.
@@
A $k$-VASS can be seen as an infinite-state labelled transition system in which each configuration is a pair $(s,u)\in Q\times\mathbb{N}^{k}$. We denote such configuration as $s(u)$. We define the norm of the configuration as the norm of $u$. A transition $t=(s,a,v,s^{\prime})\in\delta$ can be fired in a configuration $q(u)$ if and only if $q=s$ and $u^{\prime}\geq 0$ where $u^{\prime}=u+v$. After firing the transition configuration is changed to $s^{\prime}(u^{\prime})$. We also define run as a sequence of transitions, which can be fired one after another from some configuration. For a run $\rho=t_{1}t_{2}\ldots t_{n}$ we write $\textsc{eff}(\rho)$ for $\Sigma_{i=1}^{n}\textsc{eff}(t_{i})$. If we want to say something only about $j$th (for $j\in[k]$) entry of the $\textsc{eff}(\rho)$ we write $\textsc{eff}_{j}(\rho)$. For a run $\rho=t_{1}t_{2}\ldots t_{n}$ and $j\in[k]$ we define also $\textup{maxdrop}_{j}(\rho)$ as $\max_{i\in[0,n]}|\textsc{eff}_{j}(t_{1}\ldots t_{i})|$ and $\textup{maxdrop}(\rho)$ as $\max_{i\in[k]}\textup{maxdrop}_{i}(\rho)$. We also define $\textup{support}(\rho)$ as $\textup{support}(\textsc{eff}(\rho))$. We say that run $\rho$ is from configuration $q(u)$ to $p(u^{\prime})$ if the sequence of transitions can be fired from $q(u)$ and the final configuration is $p(u^{\prime})$. We say that a run is a loop if the state of its initial configuration is the same as the state of the final configuration. For two runs $\rho_{1}=\alpha_{1}\ldots\alpha_{n}$ and $\rho_{2}=\beta_{1}\ldots\beta_{k}$ if sequence $\alpha_{1}\ldots\alpha_{n}\beta_{1}\ldots\beta_{k}$ is a run $\rho$ then we can write $\rho=\rho_{1}\rho_{2}$. We say, that a run $\rho=t_{1}\ldots t_{n}$ is over $w=\lambda_{1}\ldots\lambda_{n}\in(\Sigma\cup\{\varepsilon\})^{*}$ (or reads $w$) if and only if for each $i\in[n]$ transition $t_{i}$ is over $\lambda_{i}$. We denote by $w(\rho)$ the word read by $\rho$. We say that the length of a run $\rho$ is equal to $n$ if it consists of $n$ transitions. For the length of the run, we write $|\rho|$.
@@
We say, that VASS is $\varepsilon$-free if there is no transition over $\varepsilon$. In this work, unless stated otherwise, we work with $\varepsilon$-free VASSs. Languages of VASSs. A run of a VASS is accepting if it starts in an initial configuration $c_{0}\in I$ and ends in an accepting configuration. For a VASS $V$ we define its language as the set of all words read by accepting runs and denote it as $L(V)$. We mostly consider configuration to be accepting (also final) if and only if it covers some configuration from the set of final configurations $F$. That means configuration $q(v)$ is accepting if and only if there exists a configuration $q(v^{\prime})\in F$ such that $v\geq v^{\prime}$. Languages defined this way are called coverability languages.
@@
We sometimes consider reachability languages in which run ending in configuration $q(v)$ is accepting if and only if $q(v)\in F$. Sometimes, we consider VASSs, where the set of accepting configurations is infinite, but has a specific form. For instance we consider downward-VASSs, where the set of accepting configurations is possibly infinite, but it is downward-closed. In this type of VASSs a run ending in configuration $q(v)$ if and only if $q(v)\in F$ where $F$ is a downward-closed set of accepting configurations. We say, that VASS $V$ is deterministic if it has only one initial configuration, it is $\varepsilon$-free and for each state $q$ and each letter $a\in\Sigma$ there is at most one transition over $a$ leaving the state $q$. We say, that VASS $V$ is $k$-ambiguous if and only if for every $w\in L(V)$ we have at most $k$ accepting run over $w$. We also say then, that $L(V)$ is a $k$-ambiguous language. For a $k$-VASS consider a function $r:(Q\times\mathbb{N}^{k}\times\delta)^{*}(Q\times\mathbb{N}^{k})\times\Sigma\rightarrow\delta^{*}$ that, given a history of the run (configurations and taken transitions), current configuration $q(v)$ and a next letter $\lambda\in\Sigma$, returns a sequence of transitions over $\lambda$, which can be fired from $q(v)$. Let us call $r$ a resolver. We say, that $k$-VASS $V$ is history-deterministic if and only if it has one initial configuration and there exists a resolver $r$ such that for each $w\in L(V)$ run $\rho$ over $w$ from the initial configuration given by the resolver is accepting.
@@
We denote by Det, Hist, k-Amb, BAmb and NonDet the class of languages of respectively deterministic, history-deterministic, $k$-ambiguous , all boundedly-ambiguous and fully nondeterministic VASSs languages. We sometimes call the class 1-Amb the class of unambiguous VASSs languages. Well-quasi order. Quasi-order $\preceq$ defined on set $X$ is a relation satisfying:
@@
For every $d\in\mathbb{N}$ and finite set $Q$ we have, that $\preceq$ defined on $Q\times\mathbb{N}_{\omega}^{d}$ as for each $q_{1},q_{2}\in Q$ and $v,v^{\prime}\in\mathbb{N}_{\omega}^{d}$ we have $(q_{1},v)\preceq(q_{2},v^{\prime})\iff\forall_{i\in[d]}v_{i}\leq v_{i}^{\prime}\land q_{1}=q_{2}$ is a well-quasi-order .
{toappendix}

### v2

@@ Abstract
The aim of this paper is to deliver broad understanding of a class of languages of boundedly-ambiguous VASSs, that is $k$-ambiguous VASSs for some natural $k$. These are languages of Vector Addition Systems with States with the acceptance condition defined by the set of accepting states such that each accepted word has at most $k$ accepting runs. We develop tools for proving that a given language is not accepted by any $k$-ambiguous VASS. Using them we show a few negative results: lack of some closure properties of languages of $k$-ambiguous VASSs and undecidability of the $k$-ambiguity problem, namely the question whether a given VASS language is a language of some $k$-ambiguous VASS. In fact we show an even more general undecidability result stating that for any class containing all regular languages and only $k$-ambiguous VASS languages for some $k\in\mathbb{N}$ it is undecidable whether a language of a given $1$-dimensional VASS belongs to this class. Finally, we show that the regularity problem is decidable for $k$-ambiguous VASSs.
@@
vector addition systems, Petri nets, unambiguity, bounded-ambiguity, languages
@@ 1 Introduction
Determinism is a central notion in computer science. Deterministic systems often allow for more efficient algorithms. On the other hand, usually deterministic systems are less expressive, so in many cases for a non-deterministic system there is no equivalent deterministic system and one cannot use more efficient techniques. For those reasons, there is recently a lot of research devoted to various notions restricting nondeterminism in a milder way than determinism. The hope is that systems having the considered properties are more expressive than the deterministic ones, but still allow for robust algorithms design. One prominent example of such a notion is unambiguity; a system is unambiguous if for every word there is at most one accepting run over this word. In the last decade unambiguous systems were intensely studied and for various classes of infinite-state systems the unambiguous case turns out to be much more tractable [DBLP:conf/dcfs/Colcombet15, DBLP:conf/icalp/Raskin18, DBLP:conf/stacs/MottetQ19, DBLP:conf/concur/CzerwinskiFH20, DBLP:conf/icalp/GoosK022]. Similar notions were also investigated recently, like $k$-ambiguity (each word is accepted by at most $k$ runs) and history-determinism (a weakened version of determinism), in both cases one can design more efficient algorithms in some cases [DBLP:conf/concur/CzerwinskiH22, history-deterministic].
@@
In this paper we focus on studying milder version of determinism for Vector Addition Systems with States (VASS). VASSs and related Petri nets are popular and fundamental models of concurrency with many applications both in theory and in practical modelling [DBLP:journals/siglog/Schmitz16, Section 5]. Languages of VASSs with restricted nondeterminism were already studied for several years, mostly with the acceptance condition being the set of accepting states. In [DBLP:conf/concur/CzerwinskiFH20] it was shown that the universality problem for unambiguous VASSs is decidable in ExpSpace, in contrast to Ackermann-completeness of the problem for VASSs without that restriction. In [DBLP:conf/concur/CzerwinskiH22, DBLP:journals/lmcs/CzerwinskiH25] the language equivalence problem was considered for unambiguous VASS and more generally for $k$-ambiguous VASSs and it was shown to be decidable and Ackermann-complete, in contrast to undecidability in general [DBLP:journals/tcs/ArakiK76], even in dimension one [DBLP:conf/lics/HofmanMT13, Thm. 20]. The choice of universality and equivalence problems is deliberate here. The complexity of the seemingly more natural language emptiness problem (or equivalently of the reachability problem) does not depend on the unambiguity assumption. Indeed, one can always transform a system to a deterministic one by assigning all the transitions unique labels without affecting nonemptiness. Therefore, in order to observe a difference in complexity or decidability after restricting nondeterminism, one should consider problems, which do not simply ask about the existence of some object. Some natural problems concerning languages of that type are the universality problem and the language equivalence problem and these two problems were extensively studied for many models with the unambiguity assumption or some modification of it. First of all, the universality and equivalence problems are solvable in PTime [DBLP:journals/siamcomp/StearnsH85] (and even in NC2 [DBLP:journals/ipl/Tzeng96]) for unambiguous finite automata (UFA), in contrast to PSpace-hardness of both problems for NFA. The universality problem was also shown decidable for unambiguous register automata in a sequence of papers [DBLP:conf/stacs/MottetQ19, DBLP:conf/stacs/BarloyC21, DBLP:conf/icalp/CzerwinskiMQ21] with improving complexity, which contrasts undecidability in the case without restricted nondeterminism [DBLP:journals/tocl/NevenSV04, Thm 5.1]. This line of research culminated in a very elegant contribution [DBLP:conf/lics/BojanczykKM21, DBLP:journals/theoretics/BojanczykFKM24], which showed in particular that not only the universality problem, but also the equivalence problem is solvable in ExpTime for unambiguous register automata, and in PTime in the case of fixed number of registers. Another popular restriction of nondeterminism was also studied recently: history-determinism. A system is history-deterministic if its nondeterminism can be resolved on the fly, based on the history of a particular run. The equivalence and inclusion problems were shown to be decidable for one-dimensional history-deterministic VASSs [DBLP:conf/fossacs/PrakashT23], but undecidable for two-dimensional history-deterministic VASSs [history-deterministic].
@@
Despite a lot of research on algorithms for unambiguous and boundedly-ambiguous ($k$-ambiguous for some $k\in\mathbb{N}$) VASS not much is known about the class of languages recognisable by such models. In particular, to our best knowledge, till now it was even not known whether there exist any VASS language (we consider acceptance by states), which is not a language of an unambiguous VASS. The reason behind this lack of knowledge was absence of any technique, which can show that a given language of a VASS cannot be recognised by an unambiguous VASS. The quest for such a technique is natural and the question deserves investigation. Analogous problems were considered for other models of computation. For finite automata the question trivialises, as deterministic automata recognise all the regular languages. However, already for weighted automata over a field the problems are highly nontrivial and were recently studied in-depth [DBLP:conf/lics/BellS23, DBLP:journals/corr/abs-2410-03444], in particular it is decidable whether a given weighted automaton is unambiguisable, so equivalent to some unambiguous weighted automaton [DBLP:conf/lics/BellS23]. The problem whether a given context-free language is unambiguisable is known to be undecidable since the 60s [undecidable_cfl, DBLP:journals/mst/Greibach68]. The aim of this paper is deepening the understanding of the class of languages recognisable by boundedly-ambiguous VASS and its subclasses.
@@
For any class $\mathcal{C}$ of languages containing all regular languages and contained in the class of all boundedly-ambiguous VASS languages it is undecidable to check whether the language of a given $1$-VASS accepting by states belongs to $\mathcal{C}$.
@@
It is undecidable whether the language of a given $1$-VASS accepting by states is recognisable by some
@@
  • $k$-ambiguous 11-VASS for given $k\in\mathbb{N}$.
@@
Organisation of the paper. In Section 2 we introduce preliminary notions and recall useful lemmas. Section 3 is devoted to showing the two main technical tools, namely Lemmas 3.2 and 3.20. In Section 4 we present closure properties and expressivity results for the classes of $k$-ambiguous VASSs. Next, in Section 5 we show our main result, namely Theorem 1.1. Theorem 6.1 is proved in Section 6. Finally, in Section 7 we discuss interesting future research directions.
@@ 2 Preliminaries
Basic notions. For $a,b\in\mathbb{N}$ such that $a\leq b$ we write $[a,b]$ for the set of integers $\{a,a+1,\ldots,b\}$. For $a\in\mathbb{N}$ we write $[a]$ for the set $[1,a]$. For a vector $v\in\mathbb{Z}^{d}$ we write $v_{i}$ for the $i$-th entry of $v$. For a vector $v\in\mathbb{Z}^{d}$ we write $\textup{support}(v)$ for the set $\{i\mid v_{i}>0\}$. For two vectors $u,v\in\mathbb{N}^{d}$ we write $u\leq v$ if for all $i\in[d]$ we have $u_{i}\leq v_{i}$. We also extend this order to $(\mathbb{N}\cup\{\omega\})^{d}$ where $\omega$ is bigger than any natural number. We write $\mathbb{N}_{\omega}$ for the set $\mathbb{N}\cup\{\omega\}$. Whenever we speak about the norm of vector $v\in\mathbb{Z}^{d}$ we mean $v||=\max_{1\leq i\leq d}|v_{i}|$. For a word $w$ we denote by $_{a}(w)$ the number of letters $a$ in the word $w$.
Downward-closed sets. A set $S\subseteq N^{d}$ is downward-closed if for each $u,v\in\mathbb{N}^{d}$ it holds that $u\in S$ and $v\leq u$ implies $v\in S$. For $u\in\mathbb{N}^{d}$ we write $u\downarrow$ for the set $\{v\mid v\leq u\}$. If a downward-closed set is of the form $u\downarrow$ we call it a down-atom. Observe, that a one-dimensional set $S\subseteq\mathbb{N}$ is downward-closed if either $S=\mathbb{N}$ or $S=[0,n]$ for some $n\in\mathbb{N}$. Thus we have either $S=\omega\downarrow$ in the first case or $S=n\downarrow$ in the second case. So equivalently a downward-closed set $D\subseteq\mathbb{N}^{d}$ is a down-atom if it is of a form $D=D_{1}\times\ldots\times D_{d}$, where for all $i\in[d]$ we have that $D_{i}$ is a downward-closed one dimensional set. For simplicity we write $D=(n_{1},n_{2},\ldots,n_{d})\downarrow$ if $D=n_{1}\downarrow\times n_{2}\downarrow\ldots n_{d}\downarrow$. Therefore each down-atom is of the form $u\downarrow$ where $u\in(\mathbb{N}\cup\omega)^{d}$. The following proposition will be helpful in our considerations.
Proposition 2.1 (Lemma 17 in [CzerwinskiLMMKS18], [Dickson]).
@@
Vector Addition Systems with States. A $d$-dimensional Vector Addition System with States ($d$-VASS) is a nondeterministic finite automaton with $d$ non-negative integer counters. Transitions of the VASS manipulate these counters. Formally, we define VASS $V$ as $V=(\Sigma,Q,\delta,I,F)$ where $\Sigma$ is a finite alphabet, $Q$ is a finite set of automaton states, $\delta\subseteq Q\times(\Sigma\cup\varepsilon)\times\mathbb{Z}^{d}\times Q$ is a transition relation, $I\subseteq Q\times\mathbb{N}^{d}$ is a finite set of initial configurations and $F\subseteq Q\times\mathbb{N}^{d}$ is a finite set of final configurations. For a transition $t=(s,a,v,s^{\prime})\in\delta$ we say that the transition is over $a$ or the transition reads letter $a$. We also write $\textsc{eff}(t)$ for the effect of transition, which is $v$. We define the norm of transition $t$ as the norm of $v$.
@@
A $d$-VASS can be seen as an infinite-state labelled transition system in which each configuration is a pair $(s,u)\in Q\times\mathbb{N}^{d}$. We denote such configuration as $s(u)$. We define the norm of the configuration as the norm of $u$. A transition $t=(s,a,v,s^{\prime})\in\delta$ can be fired in a configuration $q(u)$ if and only if $q=s$ and $u^{\prime}\geq 0$ where $u^{\prime}=u+v$. After firing the transition configuration is changed to $s^{\prime}(u^{\prime})$. We also define run as a sequence of transitions, which can be fired one after another from some configuration. For a run $\rho=t_{1}t_{2}\ldots t_{n}$ we write $\textsc{eff}(\rho)$ for $\Sigma_{i=1}^{n}\textsc{eff}(t_{i})$. If we want to say something only about $j$th (for $j\in[d]$) entry of the $\textsc{eff}(\rho)$ we write $\textsc{eff}_{j}(\rho)$. We also define $\textup{support}(\rho)$ as $\textup{support}(\textsc{eff}(\rho))$. We say that a run $\rho$ is from configuration $q(u)$ to $p(u^{\prime})$ if the sequence of transitions can be fired from $q(u)$ and the final configuration is $p(u^{\prime})$. We say that a run is a loop if the state of its initial configuration is the same as the state of the final configuration. For two runs $\rho_{1}=\alpha_{1}\ldots\alpha_{n}$ and $\rho_{2}=\beta_{1}\ldots\beta_{k}$ if the sequence $\alpha_{1}\ldots\alpha_{n}\beta_{1}\ldots\beta_{k}$ is a run $\rho$ then we can write $\rho=\rho_{1}\rho_{2}$. We say, that a run $\rho=t_{1}\ldots t_{n}$ is over $w=\lambda_{1}\ldots\lambda_{n}\in(\Sigma\cup\{\varepsilon\})^{*}$ (or reads $w$) if and only if for each $i\in[n]$ transition $t_{i}$ is over $\lambda_{i}$. We denote by $w(\rho)$ the word read by $\rho$. We say that the length of a run $\rho$ is equal to $n$ if it consists of $n$ transitions. For the length of a run, we write $|\rho|$.
@@
We say, that VASS is $\varepsilon$-free if there is no transition over $\varepsilon$. In this work, unless stated otherwise, we work with $\varepsilon$-free VASSs. Languages of VASSs. A run of a VASS is accepting if it starts in an initial configuration $c_{0}\in I$ and ends in an accepting configuration. A configuration is accepting if it covers some configuration from the set of final configurations $F$. In other words configuration $q(v)$ is accepting if there exists a configuration $q(v^{\prime})\in F$ such that $v\geq v^{\prime}$. For a VASS $V$ we define its language as the set of all words read by accepting runs and we denote it by $L(V)$. Languages defined in this way are called coverability languages. Sometimes we consider languages with other acceptance condition (reachability languages), in that case we indicate it clearly.
@@
We sometimes consider reachability languages in which run ending in configuration $q(v)$ is accepting if and only if $q(v)\in F$. Sometimes, we consider VASSs, where the set of accepting configurations is infinite, but has a specific form. For instance we consider downward-VASSs, where the set of accepting configurations is possibly infinite, but it is downward-closed. In this type of VASSs a run ending in configuration $q(v)$ if and only if $q(v)\in F$ where $F$ is a downward-closed set of accepting configurations. We say, that VASS $V$ is deterministic if it has only one initial configuration, it is $\varepsilon$-free and for each state $q$ and each letter $a\in\Sigma$ there is at most one transition over $a$ leaving the state $q$. We say, that VASS $V$ is $k$-ambiguous if and only if for every $w\in L(V)$ we have at most $k$ accepting runs over $w$. We also say then, that $L(V)$ is a $k$-ambiguous language. For a $d$-VASS consider a function $r:(Q\times\mathbb{N}^{d}\times\delta)^{*}\times(Q\times\mathbb{N}^{d})\times\Sigma\rightarrow\delta^{*}$ that, given a history of the run (configurations and taken transitions), current configuration $q(v)$ and a next letter $\lambda\in\Sigma$, returns a sequence of possibly many transitions. One of these transitions is over $\lambda$, all the other are over $\varepsilon$, and the sequence can be fired from the current configuration $q(v)$. Let us call $r$ a resolver. We say, that $d$-VASS $V$ is history-deterministic if and only if it has one initial configuration and there exists a resolver $r$ such that for each $w\in L(V)$ run $\rho$ over $w$ from the initial configuration given by the resolver is accepting.
@@
We denote by Det, Hist, k-Amb, BAmb and NonDet the class of languages of respectively deterministic, history-deterministic, $k$-ambiguous , all boundedly-ambiguous and fully nondeterministic VASSs languages. We sometimes denote by 1-Amb the class of unambiguous VASSs languages. By d-NonDet we denote the class of languages of $d$-VASSs.
Well-quasi order. A quasi-order $\preceq$ defined on set $X$ is a relation satisfying:
@@
For every $d\in\mathbb{N}$ and finite set $Q$ we have, that $\preceq$ defined on $Q\times\mathbb{N}_{\omega}^{d}$ as for each $q_{1},q_{2}\in Q$ and $v,v^{\prime}\in\mathbb{N}_{\omega}^{d}$ we have $(q_{1},v)\preceq(q_{2},v^{\prime})\iff v\leq v^{\prime}\land q_{1}=q_{2}$ is a well-quasi-order .